# Cognitive Visual-Language Mapper: Advancing Multimodal Comprehension with Enhanced Visual Knowledge Alignment

### v1

@@ 5 Conclusion
In this work, we introduce the Cognitive Visual-Language Mapper (CVLM), an innovative approach that goes beyond the conventional alignment of visual and textual descriptions by incorporating visual-knowledge alignment. Specifically, we have developed a Visual Knowledge Aligner (VKA) that facilitates the projection of visual knowledge by acting as a bridge between the vision encoder and LLM. Additionally, we have integrated a Fine-grained Visual Knowledge Adapter (FKA) during the multimodal instruction tuning stage, which is designed to distill more precise knowledge pertinent to images and instructions. Our experimental findings demonstrate that CVLM outperforms several prominent LMMs that lack visual knowledge alignment. Our ablation studies highlight the effectiveness of VKA and FKA.
@@
There are several limitations to our work: 1) The knowledge representations gained by VKA may be inaccurate due to the loss of visual information and errors of knowledge association. Although we introduce large-scale visual knowledge data and the FKA to enhance visual knowledge acquisition, there is still potential to improve the accuracy of visual knowledge alignment. 2) From the experimental results, we observed that the distillation vector length impacts the stability of language models infused with visual knowledge information. Hence, we still need to explore an effective and stable visual knowledge-enhanced version of CVLM, especially for its FKA component. 3) The generated content may contain some factual errors or toxic statements due to the limitations of LLMs’ generation capabilities. We also hope this work could spark further research on improving visual knowledge alignment during the construction of LMMs.

### v2

@@ 5 Conclusion
In this work, we introduce the Cognitive Visual-Language Mapper (CVLM), an innovative approach that goes beyond the conventional alignment of visual and textual descriptions by incorporating visual-knowledge alignment. Specifically, we have developed a Visual Knowledge Aligner (VKA) that facilitates the projection of visual knowledge by acting as a bridge between the vision encoder and LLM. Additionally, we have integrated a Fine-grained Visual Knowledge Adapter (FKA) during the multimodal instruction tuning stage, which is designed to distil more precise knowledge pertinent to images and instructions. Our experimental findings demonstrate that CVLM outperforms several prominent LMMs that lack visual knowledge alignment. Our ablation studies highlight the effectiveness of VKA and FKA.
6 Acknowledge
Thanks for the efforts from reviewers and action editors. This work is supported by grants: Natural Science Foundation of China (No. 62376067).
@@ 6 Acknowledge
Our work has several limitations: 1) The knowledge representations gained by VKA may be inaccurate due to the loss of visual information and errors of knowledge association. Although we introduce large-scale visual knowledge data and the FKA to enhance visual knowledge acquisition, there is still potential to improve the accuracy of visual knowledge alignment. 2) From the experimental results, we observed that the distillation vector length impacts the stability of language models infused with visual knowledge information. Hence, we still need to explore an effective and stable visual knowledge-enhanced version of CVLM, especially for its FKA component. 3) The generated content may contain some factual errors or toxic statements due to the limitations of LLMs’ generation capabilities. We also hope this work could spark further research on improving visual knowledge alignment during the construction of LMMs.
Appendix A How Wikipedia knowledge pretraining data Constructed?
The construction of our Wikipedia knowledge pretraining data leverages images and their corresponding descriptions from approximately 2 million Wikipedia pages. Each page includes multiple images, enriched with Contextual Image Captioning and section summarization to serve as associated knowledge descriptions for the images. This comprehensive dataset is derived from work[1], providing a foundational resource for our pretraining data. Data Construction Processes are:
Data Processing: The initial step involves reading and parsing the TFRecord file. To manage the vast amount of data efficiently, concurrently, a thread pool is employed to download images based on the parsed image URLs.
Filtering: It was observed that approximately 1% of the image URLs were invalid. To ensure the quality and integrity of the dataset, these invalid entries were identified and subsequently removed from the dataset.
[1] WIT: Wikipedia-Based Image Text Dataset for Multimodal Multilingual Machine Learning, 2021. https://github.com/google-research-datasets/wit/blob/main/wikiweb2m.md